Summary

Check of the optimization gain after the modification to perform two pass to TT and one pass to HO1.Also an implementation of a quadratic fitting.

Test the system for bin 3 and 4, magnitude 15.7 to 17.3. The seeing turbulence was 0.8"

Details:

The start of the system went well.

Close loop set and the OCAM viewer wnet NOT LIVE with the loop close in the secondary and commands arrving to it.
Debugging Alfio and Fabio. Test different alternativies workaround but when the loop is close the OCAM was NOT LIVE and the secondary still receving bad frames.

It was observed that the WFSframes at the WFS showed to be corrupted and it was ONLY when the Adsec was power OFF it was availabel to recover the system to the normal operations.
There is a suspicious that bad frames "corrupted" between the Adsec adn the WFS triggered a fault in the WFS masterdiagnostic. This issue is under investigation by Arcetri and a possible solution of a better handshaking protocol could mitigate this issue.

Restore the system.

System ready for check the bin3 and bin4 optimization gain.

The magnitude from 15.7 to 16.3 were tested without problems.

The new implementation of bin3 and bin4 went well and it went through all steps (including the check flux ) to optical gain optimizations. The processes didn't show any execution problems. More details from Enrico emails.
Today-> preset timings. Next to use NONINTERVENTION MODE.

Let me know if the timings are corrects.
preset timings:* the sequence was not followed after each step completion.
20:37:24.3 PresetAO mag 16.3
20:39:33 center Star
2039:56.1 CenterPupils
20:41:04 check flux ->457Hz
20:42:45.1.9 OptGain
20:43:26.9 OptGain Done.

20:43:37.6 Gopt

20:44:09.2 Done Gopt

TN#_204500

Shutter 48 mag 15.7

20:49:24.0t PresetAO mag 16.3

20:50:12.9 CenterStar

20:51:12 CenterPupil

20:52:21.49 CheckFlux ->489Hz

20:53:08.25 OptGain

20:53:48.3 done OptGain

20:53:59.77 DoneGopt

TN#205515

Shutter 43 mag 17.3

20:58:46.12 PresetAO mag 16.3

20:59:36.7 CenterStar

21:00:02.14 CenterPupil.

21:01:41.34 CheckFlux->247Hz

21:03:06.24 OptGain

21:04:26 Done OptGain

21:05:25.79 Gopt.

21:06:098.39 DoneGopt

TN#210707
  • Mag 16.3 .shutter 46 forced
    autogain 20210211_202415
    gopt
    TN: 20210211_203342 203537 203555
    Executed checkflux = R=16.6 mode to bin4 Fre=457Hz
    autogain 20210211_204241
    gopt
    TN# 20210211_204420 204440 204500
  • Mag 15.7. shutter 48
    autogain 20210211_205308
    gopt
    TN 20210211_205448 205502 205515
  • Mag 17.4 . shutter 43 -->247Hz
    autogain 20210211_210306
    gopt
    TN 20210211_210616 210645 210707

-- JuanCarlosGuerra - 12 Feb 2021
Topic revision: r1 - 12 Feb 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback